OpenStudy (anonymous):

Solve: -4(x + 1) - 3 = -3(x - 4) 1.-19 2.19 3.2 4.0

OpenStudy (anonymous):

-4x-4-3=-3x+12 =>-4x+3x=12+7 =>-x=19 =>x=-19

OpenStudy (anonymous):

First you distribute the -4 by multiplying by both terms in the parentheses, and similar to the other one. \(-4x-4-3=-3x+12\) Now simplify

OpenStudy (anonymous):

thanks